OpenStudy (anonymous):

Quadrilateral MNTU below is a parallelogram. If m∠U = 5x and m∠T = 2x – 9, what is the measure of ∠T? A. 3° B. 15° C. 27° D. 45°

OpenStudy (anonymous):



OpenStudy (anonymous):

2 (5x) + 2(2x - 9) = 360 10x + 4x - 18 = 360 14x - 18 = 360 14x = 360 +18 14x = 378 378/14 = x x = 27

OpenStudy (anonymous):

thanks for all the steps :)

OpenStudy (anonymous):

No problem :)